OpenStudy (anonymous):

Twice a number is 12 more than five times the number. Find the number.

OpenStudy (anonymous):

equation 2x=5x+12 2x-5x=12 -3x=12 x=-4

OpenStudy (anonymous):

Okay, so we had our equation wrong. Thank you!

OpenStudy (anonymous):

welcome

OpenStudy (karatechopper):

-4